OpenStudy (anonymous):

need help in algebra angle 1 (6x-32) angle 2 (2x-20) angle 3 (3y-8) angle 4 (2x-20)

OpenStudy (anonymous):

geometry not algebra

OpenStudy (anonymous):

vertical angles are what?

OpenStudy (anonymous):

congruent

OpenStudy (anonymous):

no they are equal to each other... But first, what is the question?

OpenStudy (anonymous):

Sorry, should of said "think of them as equal to each other" Congruent is correct..but better to think of them as equal

OpenStudy (anonymous):

i attatched a photo and where they go is in the problem

OpenStudy (anonymous):

correct, but what are they asking you to do? I see the picture, but is the question there also?

OpenStudy (anonymous):

Find the values of x and y in the diagram below

OpenStudy (anonymous):

ok... so lets pick any two vertical angle

OpenStudy (anonymous):

1 and 3 are vertical as well as 2 and 4, right?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

are you sure that angle 2 and 4 data given in your question right? they are the same thing

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

angle 1 (6x-32) angle 3 (3y-8) angle 2 (2x-20) angle 4 (2x-20) lets use the straight angles to find the value of X first. So you understand that straight angles total 180

OpenStudy (anonymous):

yes ma'am.

OpenStudy (anonymous):

6x - 32 + 2x -20 = 180 8x - 52 = 180 8x = 180 - 52 8x = 128 x = 16

OpenStudy (anonymous):

So, I just chose angle 1 and 2 and set their SUM to 180 - solved for x

OpenStudy (anonymous):

now lets find Y

OpenStudy (anonymous):

the only angle with a Y variable is 3; so we know that 1 and 3 are vertical. We know this because they are ACROSS from each other.

OpenStudy (anonymous):

okay im following

OpenStudy (anonymous):

remember we know the value of X now so angle 1 = 6x-32 <1 = 6(16) - 32 <1 = 64

OpenStudy (anonymous):

Good to know :)

OpenStudy (anonymous):

so now, we set <1 = <3 64 = 3y - 8 solve for Y 64 + 8 = 3Y 72 = 3Y y = 24

OpenStudy (anonymous):

what is <1 ?

OpenStudy (anonymous):

There you go..values for X and Y You are NOW sure of Vertical angles, right? Next time..just look to see which angles have same variable and see if you can apply straight line theory or vertical angle theory to solve for the variable <1 = angle 1

OpenStudy (anonymous):

once you solve for that variable...look to see which angle have the other variable you need. Plug in Number value for the first one found and then just solve for the 2nd one. You good?

OpenStudy (anonymous):

one sec let me go over it

OpenStudy (anonymous):

ok..np - take your time, brb

OpenStudy (anonymous):

in <1= 6(16) - 32..... where did you get the 6

OpenStudy (anonymous):

nvm i see thank you so much Ms.M you really helped me out here

OpenStudy (anonymous):

angle 1 (6x-32) from what you provided in the question

OpenStudy (anonymous):

Well good to know. Anytime...:) Happy studies.